Assuming that the MPC=.75 and that prices are constant, which of the following fiscal policies would eliminate a recessionary gap of $60 billion maintaining a balanced budget?
A. Decreasing government spending by $60 billion while raising taxes $80 billion
B. Increasing government spending by $60 billion while reducing taxes $80 billion
C. Increasing government spending by $60 billion while raising taxes $60 billion
D.Decreasing government spending by $60 billion while reducing taxes $60 billion
E. Increasing government spending by $60 billion while raising taxes $80 billion

If the economy is near the full-employment level of output the AD/AS analysis of a change in government spending or taxation will be different from the Keynesian analysis.
T or F
1
Expert's answer
2016-06-17T09:54:03-0400
C. Increasing government spending by $60 billion while raising taxes $60 billion
multiplicator of government spending=1/MPS=1/(1-MPC)=1/(1-0.75)=4
multiplicator of taxes=-MPC/MPS=-MPC/(1-MPC)=-0.75/(1-0.75)=-3
Increasing government spending by $60 billion will increase GDP by 60*4=240 billion while raising taxes $60 billion will decrease GDP by 60*3=180 billion.aise
As a result, GDP will raise by 60 billion.